# WiLLM: An Open Wireless LLM Communication System

### v1

@@ 1.4. Dataset
The World’s First LLM Wireless Communication Dataset and Benchmarks (Section 5). We introduce the first comprehensive LLM communication dataset with 1,649,996 records, each containing 58 communication metrics synchronized across UE-gNB-CN interfaces. This enables holistic cross-layer analysis of LLM communication patterns. In addition, we present two metrics, LAREI and LSEQ, as benchmarks that offer a standardized methodology for wireless LLM service evaluation.
1.5. Case Study
Smart Glasses Case Study (Section 6). The practical utility of WiLLM is validated through a smart glasses application that implements gesture-triggered queries and scene interpretation. Based on the principle of user-perceived continuity in human-computer interaction, we establish a response latency target of approximately 2 seconds, achieving optimal slicing through offline statistical analysis and online learning methodologies.
2. Related Work and Comparison
2.1. Wireless Communication for LLM
The convergence of LLMs with wireless systems has precipitated prolific theoretical advancements characterized by a critical implementation deficit. Despite many approaches being proposed, the field confronts a fundamental disjunction between algorithmic innovation and operational realization.
Architectural paradigms manifest in three configurations: 1. Cloud-edge frameworks (Chen et al., 2024b) (Wu et al., 2020) (Dong et al., 2024) establish a hierarchical computational distribution by allocating processing tasks between the cloud and edge devices. This approach combines the high computational power of the cloud with the low latency characteristics of the edge. For example, NetGPT (Chen et al., 2024b) relies on cloud-edge collaboration in personalized generative services. However, the dynamic nature of wireless channels limits its adaptability. Reference (Wu et al., 2020) explores the potential of cloud-edge computing, while (Dong et al., 2024) highlights the challenges of implementing edge AI generated from cloud-based LLMs. 2. Edge computing deployment (Cai et al., 2024) (Nelson and Toloui, 2023) (Tang et al., 2023) runs LLMs directly on edge servers or devices, enhancing performance through quantization techniques. For instance, Edge-LLM (Cai et al., 2024) optimizes edge adaptability using unified compression and adaptive layer voting, while (Nelson and Toloui, 2023) demonstrates hardware-supported edge deployment cases. Nevertheless, the limited capacity of edge devices remains a bottleneck, as further validated in (Tang et al., 2023). 3. Distributed architectures (Xue et al., 2024) (Mitra et al., 2024) decompose LLM computational tasks across multiple nodes to fully utilize distributed resources. For example, WDMoE (Xue et al., 2024) proposes a wireless distributed framework based on mixture-of-experts, though its validation in congested scenarios remains insufficient.
Resource optimization methodologies span cross-modal approaches (JPPO (You et al., 2024)) integrating prompt compression with power optimization; adaptive partitioning strategies (Chen et al., 2024a) reconfiguring computational boundaries under network fluctuations; and collaborative frameworks (He et al., 2024) optimizing offloading decisions across distributed resources. Service integration architectures encompass specialized slicing methodologies (LLM-Slice (Liu et al., 2024)), intelligent orchestration approaches (Dandoush et al., 2024), and long-context integration strategies (Xu et al., 2025; Zhang et al., 2025).
Bottlenecks and our solutions: The field faces three bottlenecks: theory-centric validation lacking real-world relevance, insufficient cross-layer integration, and unproven business viability. WiLLM addresses these through an empirical platform bridging theory and implementation. Unlike conventional edge server strategies that incur coordination inefficiencies, we deploy LLMs on GPU-equipped infrastructure at the backhaul-RAN nexus, enhancing computational capacity, resource optimization, and infrastructure integration.
2.2. Related Systems and Comparisons
In this section, we evaluate prominent open-source systems for their capabilities across a range of technical features essential for modern wireless networks, including OAI, srsRAN, Open5GS, FlexRIC, and Janus. While these systems contribute significantly to wireless network research and development, they exhibit limitations in fully supporting the comprehensive demands of next-generation networks. OAI provides a complete 4G/5G protocol stack implementation, spanning from the physical layer to the network layer. Its modular design supports academic research and prototyping (Nikaein et al., 2014). srsRAN is a lightweight and highly customizable 4G/5G software suite (Gomez-Miguelez et al., 2016), offering flexibility in scheduling and QoS management. Open5GS implements comprehensive 5G core network functions, including AMF, SMF, and UPF components (Barbosa et al., 2024), establishing a solid foundation for network integration. FlexRIC is a programmable RAN SDK focused on 5G multi-tenancy and low-latency optimization (Schmidt et al., 2021). Janus is a programmable RAN platform developed by Microsoft, based on the O-RAN architecture, supporting dynamic service models (Foukas et al., 2023). While these systems offer essential assistance in certain aspects of wireless networking, they exhibit inadequate support for general technical features like UE compatibility, slice flexibility, and specialized support for LLM communications, etc. Detailed objective metric comparisons are presented in Table 2 and Table 1 in the Appendix.
2.2.1. Research Gaps and Positioning of WiLLM
While existing open-source communication platforms provide valuable research infrastructure, comprehensive analysis reveals significant operability impediments that restrict their adoption beyond specialized research communities. As quantified in Table  1 and 2, current systems demonstrate substantial limitations in critical usability dimensions. The requirement for native slicing support at the UE level presents a particularly problematic constraint, effectively excluding standard commercial devices without specialized protocol modifications (Saboorian and Xiang, 2017; Andrade and Wickboldt, 2025). This limitation fundamentally contradicts the purpose of experimental testbeds, which should facilitate algorithm validation across diverse implementation scenarios rather than constraining experimental configurations to narrow protocol-specific deployments. Furthermore, these resulted in an undesirable ecosystem fragmentation where testbeds remain utilized primarily by their own development communities, significantly limiting cross-disciplinary innovation.
WiLLM aims to bridge the technological gap between wireless communication and LLM systems by extending the OAI architecture to build the first open-source wireless communication system designed specifically for LLM services.
3. Paradigm
LLMs as transformative technologies in wireless communications call for fundamental network architecture reconceptualization. Despite edge computing’s partial success in LLM deployment, its device-centric approach inherently limits scalability and optimization. We propose a novel core network paradigm that positions LLM resources at the nexus of high-capacity backhaul and wireless access networks, enabling comprehensive computational-communication resource integration that transcends traditional edge computing constraints.
3.1. Deployment Paradigm for LLM Services
Current LLM service deployment architectures exhibit fundamental architectural limitations that constrain their performance envelopes. As illustrated in Figure 2, traditional cloud-based configurations necessitate traversal of multiple network boundaries: from UE through gNB and CN to remote LLM servers via public internet infrastructure. This architectural paradigm introduces non-deterministic network latency due to Internet Gateway traversal, unpredictable congestion during peak utilization periods, and critically, a decoupling of computational and communication resource management planes that prevents cross-layer optimization.
[…]
The technical implementation comprises three principal architectural components:
  1. (1)
    Deterministic Service Path Integration: As shown in Figure 2, our implementation establishes a direct service flow from UE through gNB to the CN-GPU integrated infrastructure. This configuration eliminates the indeterministic routing and potential congestion points inherent in public internet traversal, enabling guaranteed service-level agreement enforcement through contiguous administrative control of the entire transmission path.
  2. (2)
    Hierarchical Slice-Based Resource Orchestration: The implementation leverages the slice management mechanisms detailed in Section 3.3 to establish fine-grained coordination between communication and computational resources. Rather than treating these as separate resource domains (as in conventional implementations), our architecture implements a unified resource allocation framework where slice policies govern both radio resource allocation at the gNB and computational resource distribution at the CN-GPU nexus.
  3. (3)
    End-to-End Service Flow Management: The architecture implements a comprehensive flow control mechanism that extends from UE request initiation through transmission scheduling, computational resource allocation, and response delivery. This coordinated approach enables advanced optimization techniques such as predictive resource allocation, where the system anticipates computational requirements based on request characteristics and proactively configures both network and computational resources accordingly.
3.2. Transcending Edge Computing Limitations
The core network deployment paradigm for LLM services introduces fundamental architectural, operational, and infrastructural enhancements that transcend edge computing’s intrinsic limitations. The hierarchical “Tree-Branch-Fruit” slicing framework (Section 3.3) enables differentiated, SLA-governed LLM orchestration, surpassing edge computing’s localized resource allocation. The cross-layer optimization framework (Sections 4.2) synergistically coordinates resources across network strata, realizing dynamic adaptation unattainable through edge computing’s device-centricity. Strategic LLM resource placement at the backhaul-RAN confluence forges a computational-communication symbiosis, transcending edge deployments’ latency-processing tradeoffs. Empirical insights from the WiLLM dataset (Sections 5) corroborate the core network paradigm’s transformative potential, illuminating performance benefits and optimization opportunities eclipsing edge computing. This paradigm establishes the foundation for the “Tree-Branch-Fruit” slicing architecture, facilitating a technical environment where both communication resources and computational capabilities can be allocated through unified slice policies. The implementation enables an evolutionary deployment path that preserves investment in existing infrastructure while advancing sophisticated LLM service delivery models.
3.3. “Tree-Branch-Fruit” Slicing Architecture
Figure 3. Tree-Branch-Fruit network slicing architecture with hierarchical service organization.
Figure 4. Slice function architecture of WiLLM with integrated UE, gNB, Core Network, and Edge Server Components (Liu et al., 2024).
Building upon our core network deployment paradigm, we propose a novel “Tree-Branch-Fruit” slicing architecture that transforms conventional network slicing approaches into a comprehensive service provisioning platform. This multi-tier architecture, illustrated in Figure. 3, establishes a hierarchical framework that maintains backward compatibility with existing slicing mechanisms while introducing specialized LLM service slices. In this architecture:
  • Tree (gNB): Represents the foundational radio access network infrastructure, serving as the common entry point for all network traffic and the root of the slicing hierarchy.
  • Branch Slices: Correspond to traditional communication service slices defined in 5G standards, including enhanced Mobile Broadband (eMBB), Ultra-Reliable Low-Latency Communication (URLLC), and massive Machine Type Communication (mMTC). These branch slices are not the focus and implementations of this work. These branches maintain backward compatibility with existing network functions.
  • Fruit Slices: Denote specialized LLM service slices that extend from branch slices, each configured for specific application domains and computational requirements. Figure. 3 presents illustrative examples of potential fruit slice configurations, including various LLM service use cases with different model sizes tailored to specific application requirements.
The examples illustrated in Figure. 4 demonstrate how different application domains might be served by appropriately sized language models, with computationally intensive applications connected to branches supporting higher bandwidth and more powerful models, while lightweight IoT applications utilize smaller, more efficient models. These examples are not prescriptive implementations but rather illustrative use cases demonstrating the flexibility of the proposed architecture. This architecture establishes a “slice-as-a-LLM-service” paradigm that facilitates efficient multi-user and multi-slice coordinated scheduling through cross-layer resource optimization. The hierarchical structure enables:
  • Differentiated Service Provisioning: Telecom operators can provide LLM services with varying performance guarantees and computational capacities tailored to specific application requirements and price points.
  • Resource Isolation: The architecture enforces strict isolation between service slices, preventing performance degradation in one service from affecting others and enabling guaranteed quality-of-service levels.
  • Domain-Specific Optimization: Each fruit slice can be independently optimized for its specific application domain, with tailored model selection, inference parameters, and resource allocation strategies.
  • Modular Service Evolution: New LLM services can be deployed as additional fruit slices without disrupting existing services, facilitating incremental innovation. Our architecture achieves forward compatibility via application-layer tunneling that transcends conventional slicing constraints. This method encapsulates LLM traffic within standard application protocols, enabling infrastructure compatibility while maintaining architectural autonomy. The approach provides telecommunications operators with implementation flexibility independent of underlying radio network slice capabilities. This strategic decoupling of service implementation from network infrastructure allows fruit slices to evolve asynchronously from hardware upgrade cycles or standardization timelines. Such architectural independence ensures operational resilience across heterogeneous networks, enabling deployment of novel LLM capabilities unfettered by the limitations of conventional slicing frameworks.
The monetary value icons associated with different fruit slices in Figure. 3 illustrate the differentiated pricing potential of this architecture, with higher-resource slices commanding premium pricing while lightweight services remain cost-effective for mass deployment. This economic framework creates sustainable incentives for network operators to invest in LLM-capable infrastructure.
The “Tree-Branch-Fruit” architecture represents an advancement over conventional slicing approaches by introducing application-specific service customization while maintaining backward compatibility with existing network functions. This novel paradigm establishes the foundation for our comprehensive WiLLM implementation, which will be detailed in subsequent sections.
4. The Proposed System
Figure 5. Comprehensive system architecture of WiLLM with UE, gNB, CN, and Edge Server Components.
This section presents the comprehensive design and implementation of WiLLM, our novel wireless communication system specifically optimized for LLM service delivery. Building upon the “Tree-Branch-Fruit” slicing architecture described in Section 3, we have constructed a complete system implementation that extends the OpenAirInterface framework with specialized components for LLM service provisioning. Figure. 5 illustrates the system architecture, encompassing UE, gNB, core network, and edge server components with their associated interconnections and functional modules.
4.1. System Architecture Overview
WiLLM adopts an innovative vertical integration architecture design, achieving end-to-end coordination from user applications to core networks through a hierarchical management framework. This architecture, centered on layered slice management, organically integrates four key subsystems: the CN subsystem built on the Open5GS framework, with enhanced edge computing capabilities through a dedicated Edge Server Bridge; the gNB subsystem that implements enhanced “Tree-Branch-Fruit” slice management mechanisms based on OpenAirInterface; the UE subsystem that integrates configuration management, slice control, and performance measurement function modules; and the edge server subsystem that carries various LLM deployments and executes corresponding slice policies. These subsystems are not simply stacked together but are tightly coupled through a hierarchical slice management framework, enabling precise scheduling and collaborative optimization of cross-layer resources, thereby ensuring the efficiency of the LLM communication system.
4.2. Implementation Framework
WiLLM extends OpenAirInterface with strategic innovations that balance implementation efficiency with transformative functionality. This approach leverages established communication foundations while introducing specialized components that fundamentally transform system capabilities:
4.2.1. Dynamic Slice Compatibility
As illustrated in Figure. 5, our implementation extends conventional network slicing capabilities with LLM-specific service slices while maintaining backward compatibility. The Slice Manager of gNB encompasses both Branch Slice-UE Mapping and Fruit Slice-UE Mapping components, with bidirectional data flows (indicated by the UE Update and UE State Report pathways) enabling dynamic reassignment of resources in response to changing service demands.
The implementation employs a virtualization layer that abstracts underlying slice implementation details, while the Branch Slice Policy Executor and Fruit Slice Policy Executor enforce service-specific policies at the gNB level. As shown in the left section of Figure. 5, these policy executors interact directly with the PRB Manager, creating a decision hierarchy that translates high-level service policies into physical resource allocations.
4.2.2. Universal UE Compatibility
To address the limitations of current UE implementations, which often lack native slicing capabilities, WiLLM incorporates a middleware-based slice access mechanism through the Slice Manager of UE component. The data flow illustrated in Figure. 5 demonstrates how this component interfaces with both the Configuration Manager and Hot Start Module, enabling UEs without native slicing support to access specialized LLM services through application-layer tunneling techniques that encapsulate service traffic within standard data streams.
This compatibility layer effectively bypasses the limitations of current community implementations of primary slicing, which remain incomplete and inconsistently supported across device ecosystems, thereby expanding the addressable user base for LLM services.
4.2.3. Multi-UE Multi-Slice Scheduling Capability
WiLLM implements a comprehensive multi-UE-multi-slice coordinated scheduling mechanism that operates through a two-phase process. The Scheduler module in Figure. 5 forms the nexus of this coordination, receiving inputs from multiple sources including the Buffer Manager (which provides Buffer Status information), HARQ Manager (which signals Retransmission Requirements), and the Slice Allocation of UE component.
The multi-UE-multi-slice coordination operates through a two-phase process:
  • A global allocation phase distributes available physical resource blocks across active slices according to their relative priorities and resource guarantees.
  • An intra-slice scheduling phase allocates resources to individual UEs within each slice based on service-specific metrics.
This hierarchical approach ensures fair and efficient resource utilization while maintaining service isolation, effectively enabling legacy devices to benefit from slice-specific performance guarantees without hardware modifications.
4.2.4. Dual-Mode Resource Scheduling
WiLLM implements complementary “separated” and “embedded” resource scheduling methodologies to accommodate diverse operational requirements and infrastructure constraints:
  • Embedded Scheduling Mode: The embedded approach implements slice-aware resource allocation directly within the scheduler pipeline, as formalized in Algorithm 1. Figure. 5 illustrates how this approach integrates with the Resource Grid and Spectrum Usage Tracking components to maintain awareness of physical layer constraints while implementing policy decisions.
  • Separated Scheduling Mode: The separated approach decouples slice-specific policy decisions from core scheduling logic, employing an external decision engine that solves a global optimization problem and subject to resource constraints and slice isolation requirements. The data flow in Figure. 5 shows how the Resource Update pathway enables this external engine to maintain synchronization with the core scheduling system.
The interaction between the Uplink/Downlink Allocation Request pathways and the PUCCH Allocator modules for both uplink and downlink control channels (as depicted in Figure. 5) demonstrates how the system coordinates the allocation of scarce control channel resources across multiple UEs and slices.
4.2.5. Cross-Layer API Framework
WiLLM introduces a comprehensive API architecture (detailed in Figure. 16) that facilitates vertical integration between application requirements and network-level resource allocation. As shown in Figure. 5, this framework enables multiple feedback pathways:
  1. (1)
    Performance Feedback Loop: The UE State Report, Configuration/Feedback/Buffer, and UE Update connections establish a closed-loop feedback system that continuously monitors service performance and adapts resource allocation accordingly. The bidirectional nature of these connections, clearly depicted in Figure. 5, enables adaptive behavior at multiple timescales.
  2. (2)
    Service-Specific Optimization: The LLM Service ID mechanism enables adaptation of network parameters based on computational characteristics of different LLM implementations. Figure. 5 illustrates how this information flows from the Edge Server Subsystem through the Slice Manager to influence allocation decisions.
  3. (3)
    Resource Utilization Monitoring: The Resource Usage and Slice Allocation connections provide visibility into consumption patterns, establishing a cross-layer information flow that informs both immediate scheduling decisions and longer-term policy refinements.
4.2.6. Deployment Flexibility
The system supports flexible LLM deployment configurations, with the Edge Server Large Language Model Deployment component hosting multiple LLM variants with different parameter sizes as shown in Figure. 5, and the CN-Edge Server Bridge facilitating seamless integration between computational resources and network infrastructure through standardized file and network interfaces.
4.3. The Database Module
A distinctive feature of WiLLM is its integrated data collection module (shown in the upper section of Figure. 5), which continuously refines resource allocation strategies based on operational metrics. This module creates a meta-feedback loop that transcends individual communication sessions, collecting performance data from UE’s Performance, gNB Measurement, and Server Measurement components to inform the evolution of the Customized QoS Scheduler and Model components.
The system incorporates a comprehensive Database component (illustrated in the right section of Figure. 5) that persistently stores operational metrics, enabling longitudinal analysis of system performance. This data repository serves as both an analytical resource for offline optimization and a training corpus for online learning algorithms.
4.4. System Validation Infrastructure
To validate practical performance, we deployed a fully functional testbed (Figure.14 in Appendix) integrating a complete wireless communication chain with LLM service deployment. Following the logical architecture shown in Figure. 5, the core computing infrastructure employs an Intel Core i9-14900KF processor as gNB host, USRP B210 for RF frontend, and NVIDIA RTX 4090 GPU for LLM inference, supporting multi-user, multi-slice concurrent testing scenarios. The system’s operational status can be monitored through a custom GUI (Figure. 15 in Appendix). This configuration achieves complete system integration from radio access to core network, providing an ideal environment for evaluating WiLLM in practical deployments.
5. The WiLLM Dataset
We present the first comprehensive LLM communication dataset with synchronized metrics, enabling reproducible research and systematic evaluation. Appendix F details components, acquisition, synchronization, and error descriptions.
5.1. Dataset Metrics Overview
Our dataset contains 1,649,996 records with 58 metrics spanning the LLM service delivery chain. It covers four scenarios: static UE and gNB (290,653 records), dynamic UE and static gNB (363,906 records), static UE and dynamic gNB (430,369 records), and dynamic UE and gNB (565,068 records). “Dynamic” indicates dynamically allocated network slices, while “static” indicates invariant parameters throughout measurements. The dataset includes 22 metrics at the user equipment layer, 25 at the wireless access network layer, and 18 at the edge server layer.
Figures 17 and 18 confirm stable base station power and noise levels during data collection. We implemented NTP-based distributed synchronization, calibrating all devices through a common server for microsecond-level sampling precision. By recording timestamps at both the UE and server endpoints and using latency compensation algorithms, we maintained synchronization errors within $\pm$1.0 milliseconds.
5.2. Metric Architecture
The dataset provides synchronized metrics across three architectural layers: UE Metrics (Appendix Table 4) capture end-user experience through latency, timing, resolution, modalities, and payload characteristics; gNB Metrics (Appendix Table 6) document wireless conditions via radio parameters (MCS, CQI, BLER), resource allocation, transmission performance, and slice configuration; and CN with Edge Server Metrics (Appendix Table 5) measure computational aspects through inference timing, token processing, model loading, and response generation. The dataset’s value lies in its temporal synchronization, enabling precise correlation between user experience, network conditions, and computational performance. This facilitates advanced research: developing ML models for scheduling, implementing data-driven allocation, creating cross-layer analytics, and validating theoretical models. This corpus establishes a foundation for reproducible research in wireless systems supporting LLM services.
5.3. Preliminary Findings and Conclusions
Figure 6. Component distribution and latency analysis for uplink transmissions across resolution groups. The x-axis labels R1 - R6 represent various resolution groups for image requests sent to LLM, where R1 - R6 correspond to increasing resolutions, with “n” representing the sample size per group, and the lower value of “n” indicating the resolution range for each group. The left panel shows the temporal distribution of communication components (uplink, downlink, inference), the center panel presents statistical trends with standard errors, and the right panel quantifies each component’s percentage contribution to total latency, illustrating how higher resolutions affect time allocation in the transmission pipeline.
Figure 7. Component distribution and latency analysis for downlink transmissions across resolution groups. This figure illustrates the temporal characteristics of downlink transmissions stratified by resolution groupings. Similar to Figure. 6, it provides three complementary views: component distribution, statistical latency metrics, and proportional component analysis. All legend elements in the figure remain consistent with Figure. 6, with the only distinction being the utilization of downlink samples.
This section presents preliminary findings derived from our dataset, with particular focus on communication directionality characteristics and the impact of network slicing on performance metrics.
5.3.1. Uplink vs. Downlink Characteristic Analysis
Finding 1: Our dataset reveals distinct performance profiles for uplink and downlink communication scenarios. In the uplink case (UE image request $\rightarrow$ LLM inference $\rightarrow$ text response), as visualized in Figure. 6, we observe that inference time constitutes a substantial proportion of the total processing latency (74-87% across resolution groups). The relative contribution of uplink transmission time marginally increases (11-25%) with the rise in resolution.
Finding 2: In contrast, the downlink scenarios (text request $\rightarrow$ LLM text response $\rightarrow$ image response), depicted in Figure. 7, demonstrate markedly different characteristics. Downlink transmission time dominates the overall latency budget (81-86%), with inference time contributing a significantly smaller proportion (12-17%).
Conclusions: These directional performance characteristics validate WiLLM’s “Tree-Branch-Fruit” slicing architecture and LLM deployment paradigm in several critical ways:
  1. (1)
    In the uplink scenario, the largest delay comes from inference time, followed by uplink transmission time. This finding supports our strategy of deploying LLMs with GPUs in the CN. This is achieved by reducing the inference time through distributed LLM processing at the CN and minimizing uplink transmission delays by staying close to users, thereby avoiding the need to go through the public Internet.
  2. (2)
    The pronounced asymmetry between uplink and downlink transmission characteristics empirically validates our “Fruit Slice” architectural paradigm, confirming that effective monetization of LLM services requires direction-specific slice configurations that independently optimize bandwidth and computation intensive workloads, a capability that is not achievable through conventional slicing approaches.
  3. (3)
    The pronounced resource requirement asymmetry between uplink and downlink scenarios provides compelling empirical justification for direction-specific scheduling mechanisms. These findings validate the architectural necessity of dual-mode scheduling capabilities, confirming that efficient LLM service delivery requires dynamic, direction-aware resource provisioning rather than single-direction allocation policies. This increased flexibility in provisioning is a key enabler of the slice-based monetization framework.
These insights reveal that WiLLM’s paradigm and architecture, substantiated by real-world datasets, demonstrate critical advantages over conventional systems in operational dimensions. These advancements provide a robust framework for direction-specific resource optimization that traditional wireless systems cannot achieve. We provide further analysis of the reasons for the experimental phenomenon and additional notes in the appendix F.1.
5.3.2. Network Slicing Impact Analysis
Figure 8. Component distribution and latency analysis across slice groups. In the figure, the communication resources for uplink slices in Slice 1 to Slice 3 are gradually increasing. This examines performance impacts of different slice configurations on communication component distribution, highlighting distinct latency profiles for three configurations. The middle panel statistically validates performance differences, while the rightmost panel quantifies how slice configurations shift processing time proportions among uplink, downlink, and inference operations.
Figure 9. This figure presents a temporal analysis of PRB allocation patterns across three implementation scenarios: normal traffic (top), slice-enabled traffic (middle), and slice-distinguished traffic (bottom). The x-axis represents timestamps in microseconds, while the y-axis denotes “ulsch_current_rbs”, indicating the current number of resource blocks allocated to the UE. The visualization demonstrates the effectiveness of the slicing architecture in providing differentiated resource allocation while maintaining threshold compliance, with slice-distinguished traffic showing clear separation between service classes. The area below the dashed line is labeled as “basic data” with no data transmission, while the area above corresponds to “core data”.
Figure 10. The figure illustrates a temporal analysis of PRB allocation patterns across three scenarios: normal traffic (top), slice-enabled traffic (middle), and slice-distinguished traffic (bottom). The x-axis represents timestamps in microseconds, while the y-axis denotes “ulsch_current_bytes”, indicating the current transmitted bytes of the UE.
Finding 3: The effectiveness of our “Tree-Branch-Fruit” slicing architecture is demonstrated through comparative analysis of network performance under different slicing configurations. Figure. 8 illustrates how slice configuration affects the temporal distribution of processing components, with notable variations in inference time allocation (43.1% to 59.6%) and uplink transmission time (28.9% to 54.7%) across different slicing arrangements. This demonstrates that slicing indeed affects the latency of LLM communication services and its components, which supports WiLLM’s paradigm.
Finding 4: Examining PRB allocation patterns in Figure. 9 alongside actual byte transmission data in Figure. 10 reveals critical insights into slicing efficiency. Our implementation successfully achieves differentiated PRB allocation conforming to slice policies. Moreover, there exists a discrepancy between theoretical resource allocation and actual transmission patterns. Specifically, byte traffic does not scale proportionally with PRB allocation, suggesting potential resource utilization inefficiencies within the slicing framework. This indicates that due to the complex interplay of elements affecting communication, system execution mechanisms, and varying UE requests, PRB and actual communication bytes are not strictly linearly correlated.
Conclusions: These empirical observations demonstrate the critical value of WiLLM in several dimensions:
  1. (1)
    The significant variability in component allocation across slice configurations validates that WiLLM’s “Tree-Branch-Fruit” architecture enables effective LLM-specific service slicing, supporting our proposed diverse pricing patterns and underscoring the necessity of finer-grained resource allocation mechanisms.
  2. (2)
    The observed allocation-utilization gap reveals that theoretical and simulation-based approaches often fail to capture real-world complexities, confirming WiLLM’s role as an essential empirical platform bridging theory and implementation.
  3. (3)
    The non-linear relationship between resource allocation and actual transmission performance underscores the necessity of WiLLM’s cross-layer API framework and dual-mode scheduling capabilities, which enable adaptive optimization based on observed rather than simulated performance patterns.
The analysis of the causes of the observed phenomenon is presented in the Appendix F.2.
These findings substantiate WiLLM’s fundamental contribution: while traditional approaches rely on simplified resource allocation models, our system provides the empirical foundation and architectural flexibility required for truly efficient LLM service delivery in wireless environments. This empirical grounding enables telecommunications operators to achieve differentiated service quality and optimal resource utilization, a critical requirement for commercially viable LLM wireless services.
5.4. Benchmark
Analyzing datasets reveals intricate dependencies among communication resources, computational capacity, and service quality, requiring specialized metrics. As traditional network metrics fail to capture LLM service delivery’s multidimensional nature, we propose two complementary metrics for evaluating LLM wireless communication systems. The LLM-Aware Resource Efficiency Index (LAREI) quantifies the efficiency relationship between request complexity, computational demands, and resource utilization:
[…]

### v2

@@ 5.2. Dataset
The WiLLM dataset represents the first comprehensive empirical foundation specifically designed for wireless LLM research, providing synchronized multi-layer measurements essential for algorithm development and validation.
Dataset Architecture and Coverage: Our dataset comprises 1,649,996 synchronized records collected across 58 metrics spanning three system layers. The measurements cover four distinct operational scenarios with varying dynamics: static UE with static gNB (290,653 records), dynamic UE with static gNB (363,906 records), static UE with dynamic gNB (430,369 records), and dynamic UE with dynamic gNB (565,068 records). This systematic variation enables researchers to analyze system behavior under different mobility and resource management conditions.
Multi-Layer Metric Synchronization: The dataset captures 22 UE-layer metrics including request latency, response timing, image resolution parameters (320×240 to 640×480 pixels), response word count (50-200 words), and payload characteristics. At the gNB layer, 25 metrics document radio conditions (MCS, CQI, BLER, SNR), PRB allocation patterns, slice configurations, and scheduling decisions. The Core/Edge layer contributes 18 metrics measuring LLM inference timing, token generation rates, model selection (LLaVA, Llama 3.2), and GPU utilization patterns. Through NTP-based synchronization maintaining ±1.0ms precision, these metrics enable accurate correlation between user experience and system-level events.
Dataset Uniqueness: Unlike traditional wireless datasets focused on channel measurements or throughput statistics, WiLLM’s dataset explicitly captures the interplay between token generation patterns, wireless resource allocation, and user-perceived performance. The synchronized multi-layer nature reveals complex interactions impossible to observe through isolated measurements. The temporal synchronization enables researchers to correlate LLM inference events with network resource allocation decisions, exposing the dynamic bottleneck migration between computation and communication that characterizes LLM services.
Reproducibility and Standardization: The dataset includes comprehensive documentation with detailed metric definitions (the online Table 4) and collection parameters, enabling proper utilization and result reproduction. By providing standardized measurement data, researchers can evaluate their algorithms against the same realistic conditions, ensuring fair comparison across different approaches. This standardization is crucial for advancing the field beyond isolated demonstrations toward systematic performance improvements.
5.3. Case Study: Smart Glasses Application
To demonstrate WiLLM’s practical utility for application development, we implemented a smart glasses system that validates our platform’s ability to support real-world LLM applications while addressing human-computer interaction requirements.
5.3.1. Application Architecture and Integration
Figure 11. Smart glasses integrated with WiLLM. The system consists of customized smart glasses paired with a compact processing unit housing the 5G module, demonstrating practical deployment of our Tree-Branch-Fruit architecture.
Our smart glasses implementation, illustrated in Figure 11, represents a quintessential validation of WiLLM’s design principles. The hardware comprises custom-built smart glasses with an integrated camera positioned at the glabella for natural field-of-view capture, paired with a compact processing unit containing a 5G module and minimal computational resources. This configuration validates our architectural principle that resource-constrained devices can access sophisticated LLM capabilities through core network deployment.
The system architecture follows WiLLM’s design precisely: glasses function as intelligent sensors and display interfaces while computational processing occurs at GPU-equipped core network infrastructure. This separation validates our decision to position LLM resources at network convergence points rather than attempting resource-intensive edge deployment.
Figure 12. Gesture-triggered interaction paradigm demonstrating seamless integration between physical actions and LLM service delivery through WiLLM’s bidirectional communication capabilities.
The interaction paradigm, shown in Figure 12, demonstrates practical implementation of our bidirectional communication architecture. Users trigger scene analysis through intuitive gestures—extending five fingers followed by grasping motion—which captures visual frames and initiates LLM queries. This workflow validates WiLLM’s capability to handle the asymmetric traffic patterns identified in Section 2: high-resolution images flow uplink for scene understanding while generated descriptions return downlink for display.
5.3.2. Human-Computer Interaction Optimization
Performance Objectives: Human-computer interaction research establishes that user satisfaction depends more on predictable response patterns than absolute minimal latency  (Jacobs et al., 1997; Nabiyouni et al., 2017). This principle guided our optimization target: achieving stable 2-second response latency rather than pursuing minimal but variable delays. This objective balances perceptual continuity, technical feasibility, and system reliability within realistic network conditions.
Optimization Methodologies: We employed complementary approaches demonstrating WiLLM’s versatility for algorithm development:
Offline Analysis: Leveraging WiLLM’s comprehensive data collection capabilities, we systematically evaluated six slice configurations across extended operational periods. Each configuration was tested with diverse visual scenes and varying network conditions to build statistical performance models. Analysis revealed that intermediate slice allocations provided optimal balance between resource efficiency and performance consistency.
Online Learning: We implemented an Upper Confidence Bound (UCB) algorithm that continuously optimizes slice selection based on observed performance. The algorithm initially explores different configurations, gradually converging toward those consistently meeting latency targets.
Figure 13. Online learning progress for slice selection optimization, demonstrating WiLLM’s support for adaptive algorithm development and validation.
As shown in Figure 13, the online approach adapts to non-stationary conditions as network load varies or LLM model updates change computational requirements. This capability validates WiLLM’s cross-layer APIs that provide real-time feedback enabling dynamic optimization.
6. Related Works
6.1. Wireless Systems for AI/ML Services
Recent research has explored various approaches to supporting AI/ML workloads over wireless networks. Edge computing frameworks (Chen et al., 2024; Wu et al., 2020) position computational resources at network edges to reduce latency, but face fundamental limitations in supporting large language models due to power and thermal constraints. Cloud-edge collaborative architectures (Dong et al., 2024) attempt to balance computation between edge and cloud, yet struggle with the tight coupling between LLM inference and communication that our measurements reveal.
Distributed inference systems (Xue et al., 2024; Mitra et al., 2024) partition LLM computation across multiple nodes. While these approaches address computational scalability, they introduce complex coordination challenges and remain largely theoretical without practical wireless implementations. In contrast, WiLLM provides a complete, operational system that researchers can immediately use for experimentation.
Network slicing for AI services has received attention (Liu et al., 2024; Dandoush et al., 2024), but existing work focuses on simulation-based evaluation rather than real implementations. These studies identify the potential of slicing for service differentiation but lack the dynamic adaptation mechanisms necessary for LLM’s variable workloads. WiLLM’s Tree-Branch-Fruit architecture extends beyond static slicing to provide runtime adaptation based on observed performance.
6.2. Open-Source Wireless Platforms
Several open-source platforms provide wireless network implementations, each with distinct limitations for LLM service research. OpenAirInterface (Nikaein et al., 2014) offers comprehensive protocol implementations but requires deep wireless expertise and lacks LLM-specific features. srsRAN (Gomez-Miguelez et al., 2016) provides a lightweight alternative but similarly focuses on traditional communication services. Open5GS (Barbosa et al., 2024) implements core network functions without radio access capabilities, limiting end-to-end experimentation. Recent platforms like FlexRIC (Schmidt et al., 2021) and Janus (Foukas et al., 2023) introduce programmability for network intelligence but remain focused on traditional metrics and lack the cross-layer visibility essential for LLM services. Critically, all existing platforms require specialized UE hardware or protocol modifications for advanced features like network slicing, creating barriers for AI researchers. WiLLM addresses these limitations through our universal compatibility approach and comprehensive APIs that abstract wireless complexity. By building on OAI’s solid foundation while adding LLM-specific innovations, we provide a platform accessible to researchers without telecommunications expertise.
[…]